# Relativistic Leaky Surface Plasmon Accelerator in Nanostructured Carbon Nanotube Forest

### v1

@@ Abstract
In quantum-degenerate metallic carbon nanotubes (CNTs), the conduction electron gas forms a solid-state plasma environment where collective dynamics are dominant. This dense plasma is capable of supporting the excitation of surface plasmon on the CNT surface. In this study, we demonstrate a new class of microscale relativistic surface plasmons (RSPs) that, independent of laser polarisation and matching medium, can be directly excited by paraxial propagation of a high-intensity optical laser pulse through a cylindrical vacuum channel nanostructured in a CNT forest. The leaky field of the resonant RSP mode can support a non-evanescent longitudinal component with amplitude up to several $\mathrm{T}\mathrm{V}\mathrm{/}\mathrm{m}$-levels and a phase-matched focusing field. This field enables the sustained acceleration of relativistic electrons and positrons over long distances. We also investigate two possible mechanisms for electron injection, including edge injection and self-injection, both of which are feasible with current laser facilities, paving the way for experimental realisation. As well as highlighting a promising avenue for the development of ultra-compact, high-energy solid-state plasma particle accelerators, this work also opens up the potential for high-energy plasmonics.
@@ I Introduction
Plasmon is a quantum quasi-particle that represents the self-sustained collective oscillation of degenerate electrons in conductive materials, such as metals and semiconductors, driven by the long-range Coulomb interactions between conduction electrons. Plasmons can be excited by external sources, such as photons or charged particle beams [1, 2]. Within the bulk of plasma, this excitation is referred to as a bulk plasmon, characterised by the quantum energy $\hbar\omega_{p}$, where $\hbar$ is the reduced Planck’s constant and $\omega_{p}$ is the plasmon frequency, given by $\omega_{p}=\sqrt{4\pi e^{2}n_{e}/m_{e}}$ with $n_{e}$ the ambient plasma electron density and $m_{e}$ the rest electron mass. Plasmons can also be excited at the interface between a metal and a surrounding medium, forming surface plasmons (SPs). These SPs propagate along the interface with a phase velocity determined by the driving source. They exhibit unique eigenfrequencies and polarisation fields that depend on the surface geometry. The strongest excitation of SPs occurs at the resonance frequency, which for a perfectly flat metal surface is $\omega_{r}=\omega_{p}/\sqrt{1+\epsilon}$ [3, 4, 5], where $\epsilon$ is the dielectric constant of the bounding medium. For a vacuum boundary, where $\epsilon=1$, this simplifies to $\omega_{r}=\omega_{p}/\sqrt{2}$ [6]. SPs are of particular interest due to their ability to propagate over significant distances along the surface while maintaining strong field confinement in a narrow region near the boundary, enabling applications in fundamental physics [7, 8], electrochemistry [9], biosensing [10], plasmonic chip technology [11], and advanced microscopy [12, 13].
The confined strong electromagnetic (EM) fields of SPs are suitable for particle acceleration [14, 15]. In the non-relativistic plasmonic schemes, low-intensity laser pulses ($I<10^{18}\leavevmode\nobreak\ $\mathrm{W}\mathrm{/}\mathrm{c}\mathrm{m}^{2}$$) interact with sub-wavelength nanostructured targets, such as gratings, to generate strong localised fields [16, 17]. In the high-field regime driven by high-intensity laser pulses, the dynamics of plasma electrons and the behaviour of SPs become nonlinear and relativistic [18, 19]. These relativistic surface plasmons (RSPs) can sustain significantly higher field strengths than that of non-relativistic SPs and thus provides new opportunities for manipulating intense light and particle beams. For example, RSP excitation on solid targets has been shown to enhance the acceleration of protons [20] and electrons [21, 22, 23], and it plays a pivotal role in high-order surface harmonic generation [18, 24]. Recent advances in nanostructured materials [25, 26, 27, 28] have aroused significant interest in using RSPs for high-energy particle acceleration [29, 30, 31, 32]. However, the EM field of RSPs on planar surfaces decays exponentially in the normal direction, limiting their effective range and posing challenges for efficient relativistic particle acceleration. Moreover, significant challenges remain in the electrical, thermodynamic and mechanical properties of conventional metals used in these applications. Exciting RSPs on conventional metal surfaces presents several challenges under strong field conditions. Traditional plasmonic metals often cannot withstand high-intensity fields without degradation. Significant ohmic (resistive) losses result in rapid attenuation of surface plasmon waves, while excessive heating can alter the metal’s properties, induce structural damage, and compromise the stability of plasmonic modes. Additionally, nonlinear optical effects in metals, such as harmonic generation and self-phase modulation, can further complicate the control and predictability of plasmonic excitations. Addressing these limitations is critical to fully realising the potential of RSP-based systems in high-energy and high-field applications.
@@
Carbon nanotubes (CNTs) are graphene-based synthetic nanostructures formed by rolling single or multiple graphene sheets into cylindrical shapes. With inner diameters typically spanning a few nanometers and lengths extending up to millimetres, CNTs exhibit exceptional mechanical, thermodynamic, and electronic properties [33, 34]. Depending on their chirality, CNTs can behave as metallic conductors (armchair type) and act as polarisable media in a quantum-degenerate state [35, 36]. In this state, external sources, such as laser fields, can displace conduction electrons, creating electric dipoles through ponderomotive scattering. This displacement enables the excitation of SPs on the CNT surface [37]. Each carbon atom contributes approximately one conduction electron due to $sp^{2}$ hybridisation, leading to an electron number density up to the order of $10^{24}\leavevmode\nobreak\ $\mathrm{c}\mathrm{m}^{-3}$$ in individual CNTs. Unlike traditional metals, CNTs exhibit collisionless electron transport due to their unique structural, electronic, and quantum properties. This allows electrons to transport micrometre ($\mu\mathrm{m}$)-scale distances without significant energy loss [38, 39]. The resulting dense electron plasma can theoretically support extremely high acceleration gradients, reaching several $\mathrm{T}\mathrm{V}\mathrm{/}\mathrm{m}$. The hollow, vacuum-like channels within CNTs enable photon or particle beams to propagate over extended distances, generating wakefields for particle acceleration when driven by short, high-density particle beams [40] or X-ray laser pulses [27]. However, a single CNT currently cannot support efficient particle acceleration due to the mismatch between the $\mu\mathrm{m}$-scale dimensions of available photon or particle beams and the natural nanoscale dimensions of CNTs. As a result, direct experimental demonstrations are still lacking, as achieving resonant coupling between CNT structures and high-energy beams remains a challenge.
@@
A nanostructured CNT forest is an organised array of vertically aligned carbon nanotubes, resembling a dense forest [41, 42]. For instance, Fig.1 illustrates a porous configuration of such a forest where a $\mu\mathrm{m}$-scale central vacuum channel allows the stable propagation of an optical laser pulse. These structures are typically synthesised using advanced techniques like chemical vapour deposition (CVD) [43, 44]. Unlike unstructured solids, CNT forests offer significant flexibility in terms of geometry, density, and surface structure [45, 46, 47]. The density of a CNT forest can be precisely controlled by adjusting parameters such as the packing density or the volume fraction of CNTs [48], achieving typical densities in the range of $n_{e}\sim 10^{19-23}\leavevmode\nobreak\ $\mathrm{c}\mathrm{m}^{-3}$$. This density flexibility provides the RSPs excitation with exceptional tunability. For example, the characteristics of RSPs can be adjusted to exhibit predominantly electrostatic or electromagnetic behaviour [49], thereby enabling precise control over light-CNT interactions and optimised performance in high-field plasmonic and photonic applications. The dense plasma within the solid-state CNT forest supports ultra-high acceleration gradients, ranging from hundreds of $\mathrm{G}\mathrm{V}\mathrm{/}\mathrm{m}$ to tens of $\mathrm{T}\mathrm{V}\mathrm{/}\mathrm{m}$, while also providing a structurally robust medium for particle acceleration compared to traditional gaseous plasmas.
In this paper, we demonstrate that independent of laser polarisation or the matching medium, RSPs can be directly excited by a $\mathrm{T}\mathrm{W}$ laser pulse and effectively confined over a sufficiently long distance along the cylindrical surface of a quantum-degenerate CNT forest. The longitudinal component of the resulting leaky field is $\mu\mathrm{m}$-scale, non-evanescent. Due to the relativistic effect, it has a phase velocity close to the speed of light. Its amplitude can reach up to $\mathrm{T}\mathrm{V}\mathrm{/}\mathrm{m}$-level, enabling highly efficient acceleration of relativistic electrons and positrons over hundreds of $\mu\mathrm{m}$, supported by a strong, phase-matched transverse focusing field. We introduce two electron injection mechanisms: edge injection and self-injection. In edge injection, a well-defined electron beam with a narrow energy spectrum, sub-$\mu\mathrm{m}$ beam size, $\mathrm{f}\mathrm{s}$ duration, and $\mathrm{p}\mathrm{C}$ charge is trapped during the laser pulse entering the forest channel. In contrast, self-injection is a continuous process in which the emission of wall electrons into the central vacuum region leads to the self-trapping of an electron beam with a broader energy spectrum and $\mathrm{n}\mathrm{C}$ charge. Additionally, we provide a theoretical analysis of the fundamental physical mechanisms behind leaky RSP excitation and evaluate the experimental feasibility. This work offers valuable insights into advanced particle acceleration and presents promising applications for high-energy plasmonics and photonics.
II Relativistic surface plasmon and leaky field in the nanostructured CNT forest
Direct excitation of SPs at a flat interface is generally impossible due to the momentum mismatch between the incident EM wave and the SP mode, except in leaky mode configurations [50]. Here, we can show that the RSPs can be directly excited on a surface of low electron density $n_{e}<n_{c}$ by a high-intensity laser pulse, where $n_{c}=m_{e}\omega_{L}^{2}/4\pi e^{2}$ is the critical density of the laser pulse. $\omega_{L}=2\pi/\lambda_{L}$ is the laser frequency and $\lambda_{L}$ is the laser wavelength. The field energy is not fully confined to the surface but can leak into the vacuum region. However, the resulting leaky field is evanescent on a flat surface, limiting its suitability for relativistic particle acceleration, as discussed in Sec. IV.2. In contrast, a nanostructured CNT forest, with its cylindrical surface geometry, can significantly modify the spatial and temporal characteristics of the leaky field, as shown in Fig. 2.

### v2

# Leaky surface plasmon-based wakefield acceleration in nanostructured carbon nanotubes

@@ Abstract
Metallic carbon nanotubes (CNTs) can provide ultra-dense, homogeneous plasma capable of sustaining resonant plasma waves—known as plasmons—with ultra-high field amplitudes. These waves can be efficiently driven by either high-intensity laser pulses or high-density relativistic charged particle beams. In this study, we use numerical simulations to propose that electrons and positrons can be accelerated in wakefields generated by the leaky electromagnetic field of surface plasmons. These plasmons are excited when a high-intensity optical laser pulse propagates paraxially through a cylindrical vacuum channel structured within a CNT forest. The wakefield is stably sustained by a non-evanescent longitudinal field with $\mathrm{T}\mathrm{V}\mathrm{/}\mathrm{m}$-level amplitudes. This mechanism differs significantly from the plasma wakefield generation in uniform gaseous plasmas. Traveling at the speed of light in vacuum, with phase-matched focusing fields, the wakefield acceleration is highly efficient for both electron and positron beams. We also examine two potential electron injection mechanisms: edge injection and self-injection. Both mechanisms are feasible with current laser facilities, paving the way for experimental realization. Beyond presenting a promising pathway toward ultra-compact, high-energy solid-state plasma particle accelerators, this work also expands the potential of high-energy plasmonics.
@@ I Introduction
Laser plasma wakefield acceleration (LWFA) is a leading next-generation accelerator technology [1]. It enables the development of compact, portable systems (e.g., for medical and industrial applications) and large-scale scientific facilities, such as those advancing high-energy physics by extending the energy frontier [2]. The exceptional performance of LWFAs arises from their ability to generate high acceleration gradients which depends on the plasma density $n_{e}$ as $E=m_{e}c\omega_{p}/e\simeq 9.6\sqrt{n_{e}[10^{22}$\mathrm{c}\mathrm{m}^{-3}$]}\leavevmode\nobreak\ [$\mathrm{T}\mathrm{V}\mathrm{/}\mathrm{m}$]\propto n_{e}^{1/2}$, where $n_{e}$ is the ambient plasma electron density, $m_{e}$ and $c$ are rest electron mass and speed of light in vacuum respectively, and $\omega_{p}=\sqrt{4\pi e^{2}n_{e}/m_{e}}$ is the plasma frequency [3]. The recent progresses in LWFA are mainly based on the gaseous plasma where the density is typically in the range of $10^{16}-10^{18}\leavevmode\nobreak\ $\mathrm{c}\mathrm{m}^{-3}$$. This density can, in principle, support an acceleration gradient of $1-100\leavevmode\nobreak\ $\mathrm{G}\mathrm{V}\mathrm{/}\mathrm{m}$$. For example, the experiments at BELLA have accelerated electrons to $10\leavevmode\nobreak\ $\mathrm{G}\mathrm{e}\mathrm{V}$$ in $30\leavevmode\nobreak\ $\mathrm{c}\mathrm{m}$$ long plasma [4], which indicates the acceleration gradient of $30\leavevmode\nobreak\ $\mathrm{G}\mathrm{e}\mathrm{V}\mathrm{/}\mathrm{m}$$. This value scales with the theoretical prediction with the plasma density used in the experiment of $10^{17}\leavevmode\nobreak\ $\mathrm{c}\mathrm{m}^{-3}$$. To maximize LWFA capabilities—whether for miniaturizing accelerators or achieving higher particle energies—we must prioritize the use of denser plasmas. High-density gaseous plasma can be produced by high-pressure gas jets where the plasma density can reach near-critical values for optical lasers, approximately $10^{21}\leavevmode\nobreak\ \mathrm{cm}^{-3}$, and is intrinsically homogeneous [5]. However, using such high-density gas targets for plasma wakefield excitation is unfeasible because optical laser pulses would be severely attenuated in the extended density gradients at the gas jet edges. Furthermore, at densities approaching the critical density, optical lasers cannot efficiently propagate due to strong reflection and absorption, preventing effective wakefield formation.
Naturally, super-dense plasma exists in solids. Over other materials including amorphous or polycrystalline solids, crystals have long-range ordered atomic structure (periodic lattice) and then provide several advantages to be the ideal candidates for the high-density plasma wakefield excitation. Their unique crystalline structure and ionization efficiency ensure that crystals can provide higher plasma density of intrinsically homogeneous distribution, up to $10^{24}\leavevmode\nobreak\ $\mathrm{c}\mathrm{m}$^{-3}$ typically or even higher to $10^{28}\leavevmode\nobreak\ $\mathrm{c}\mathrm{m}$^{-3}$ for ultra-dense materials, e.g. diamond anvil cells. As a result, coherent behaviors, such as, resonant plasma waves (so-called plasmons), can dominate [6]. These plasmons can be excited either within the bulk of solid plasma or at the interface between a metal and a surrounding medium and are referred as bulk plasmon or surface plasmon (SP), respectively. Such solid plasma can, in principle, support ultra-high acceleration gradient, e.g. up to $\mathrm{P}\mathrm{e}\mathrm{V}\mathrm{/}\mathrm{m}$. However, natural crystal-based channel wakefield accelerator is unrealizable at current stage due to the limitation of angstrom-size channels [7, 8].
@@
It was suggested that carbon nanotubes (CNTs) may provide a solution to increase the channel size to $\mathrm{n}\mathrm{m}$ scale by compromising the acceleration gradient down to $\mathrm{T}\mathrm{e}\mathrm{V}\mathrm{/}\mathrm{m}$ [9, 10]. CNTs are graphene-based synthetic nanostructures formed by rolling single or multiple graphene sheets into cylindrical shapes. With inner diameters typically spanning a few nanometers and lengths extending up to millimeters, CNTs exhibit exceptional mechanical, thermodynamic, and electronic properties [11, 12]. Depending on their chirality, CNTs can behave as metallic conductors (armchair type) [13, 14] with the potential to support an electron number density up to the order of $10^{24}\leavevmode\nobreak\ $\mathrm{c}\mathrm{m}^{-3}$$ in individual CNTs. Hereafter, we refer to this type of metallic CNT simply as CNT for clarity. However, a single CNT currently cannot support efficient particle acceleration due to the mismatch between the $\mu\mathrm{m}$-scale dimensions of available photon or particle beams and the natural nanoscale dimensions of CNTs. As a result, direct experimental demonstrations are still lacking, as achieving resonant coupling between CNT structures and high-energy beams remains a challenge.
[…]
@@
In this paper, we propose laser-driven plasma wakefield excitation for both electron and positron acceleration in a nanostructured CNT forest channel. It is shown that the unique characteristics of this kind of plasma wakefield—including its excitation mechanisms and field dynamics—are closely linked to SPs excitation and are systematically investigated through theoretical modelling and numerical simulations. Our results show that, at moderate intensity, the CNT electrons remain confined to the inner walls of the forest channel. The electromagnetic (EM) field of resulting SPs leaks into the vacuum channel, forming a wakefield that may be used to accelerate both electrons and positrons. As the laser intensity increases, the electrons on the CNT surfaces undergo stronger oscillations, with some electrons escaping into the vacuum channel, where they become trapped and accelerated by the wakefield. In this regime, the wakefield amplitudes reach several teravolts per metre ($\mathrm{T}\mathrm{V}\mathrm{/}\mathrm{m}$). At even higher intensities, CNT electrons can freely traverse the surface, forming a plasma bubble of the high amplitude electric field, similar to those observed in gaseous plasma [24]. Two injection mechanisms are observed: edge injection and self-injection. In edge injection, a well-defined electron beam with a narrow energy spectrum (sub-$\mu\mathrm{m}$ beam size, $\mathrm{f}\mathrm{s}$ duration and $\mathrm{n}\mathrm{C}$ charge) is trapped as the laser pulse enters the forest channel. In contrast, self-injection is a continuous process in which electrons emitted from the CNT walls enter the central vacuum region, leading to the formation of an electron beam with a broader energy spectrum and $\mathrm{n}\mathrm{C}$ charge. This work provides valuable insights into the wakefield excitation in CNTs, thereby advancing the prospects for compact, highly efficient particle accelerators with experimental feasibility.
II Surface plasmon and wakefield generation in structured CNT forest channels
A high-intensity laser pulse propagating along a plasma-vacuum boundary excites the confined surface waves which can be treated as the polarisation of the plasma in response to the laser field [25]. The dispersion of SP with relativistic effects is given by
$1+\varepsilon\sqrt{\frac{n^{2}-1}{n^{2}-\varepsilon}}=0\mathrm{,}$ (1)
where $\varepsilon=1-\omega_{p}^{2}/\gamma_{e}\omega_{sp}^{2}$ and $\gamma_{e}$ is the Lorentz factor of the electron. $n=ck_{z}/\omega_{sp}=c/v_{\text{ph}}$ with $\omega_{sp}$ and $v_{\text{ph}}$ the frequency and phase velocity of the SP, respectively. $k_{z}$ is the longitudinal component of the wavevector. $\varepsilon=-1$ in resonance excitation with the frequency $\omega_{sp}=\omega_{p}/\sqrt{2\gamma_{e}}$ [26]. Eq. (1) can only be satisfied if $n>1$. $n>1$ indicates that the phase velocity $v_{ph}$ is less than $c$, leading to the dephasing problem. The phase velocity with resonance condition is given as
where $\alpha=2\gamma_{e}-1$. In the non-relativistic case $\gamma_{e}\simeq 1$, $v_{\text{ph}}\simeq 0$, indicating that SP is a standing wave with a dominant electrostatic component [25] and is unsuitable for relativistic particle acceleration. In the ultrarelativistic case $\gamma_{e}\gg 1$, for example, driven by a high-intensity laser pulse, $v_{\text{ph}}$ approaches $c$, which allows the acceleration of relativistic particles. SP excitation also depends on the surface geometry.
II.1 On flat surface
We start from the SP excitation on a solid flat surface in order to present the results with the cylindrical geometry. A typical SP excitation on the flat surface is shown Fig. 2 from a full 3D Particle-In-Cell (PIC) simulation using the code WarpX [27]. We can assume the surface (grey colour) is made of CNT forest and is initially modelled with neutral uniformly distributed Carbon atoms with the inization energies modified with the CNT properties, including weak $\pi$-bond and C-C bond ($\sigma$-bond) energies. The thermionic emission of electrons is negligible by assuming the room temperature. Field ionisation is implemented using the Ammosov-Delone-Krainov (ADK) method  [8]. The surface is positioned at $x=-3.0\leavevmode\nobreak\ $\mu\mathrm{m}$$ and extends infinitely in $x\leq-3.0\leavevmode\nobreak\ $\mu\mathrm{m}$$ and $y$ direction, and in range $0\leavevmode\nobreak\ $\mu\mathrm{m}$\leq z\leq 60\leavevmode\nobreak\ $\mu\mathrm{m}$$. Two vacuum sections of $20\leavevmode\nobreak\ $\mu\mathrm{m}$$ and $10\leavevmode\nobreak\ $\mu\mathrm{m}$$ are placed longitudinally at the head and tail of the target, respectively, for laser initialisation and electron beam diagnostics. A $5\leavevmode\nobreak\ $\mathrm{T}\mathrm{W}$$ Gaussian laser pulse of wavelength $\lambda_{L}=0.8\leavevmode\nobreak\ $\mu\mathrm{m}$$ is linearly polarised in the $y$ direction and propagates in the $z$ direction. It is focused at $z=0\leavevmode\nobreak\ $\mu\mathrm{m}$$ with root-mean-square (RMS) waist $w_{0}=3.0\leavevmode\nobreak\ $\mu\mathrm{m}$$ and duration $\sigma_{\tau}=6.0\leavevmode\nobreak\ $\mathrm{f}\mathrm{s}$$. These parameters give a normalised laser strength $a_{0}=4.0$ or intensity $I=3.4\times 10^{19}\leavevmode\nobreak\ $\mathrm{W}\mathrm{/}\mathrm{c}\mathrm{m}^{2}$$, with the field intensity of $1.46\times 10^{19}\leavevmode\nobreak\ $\mathrm{W}\mathrm{/}\mathrm{c}\mathrm{m}^{2}$$ on the surface at $x=-3.0\leavevmode\nobreak\ $\mu\mathrm{m}$$ which drives the relativistic electron dynamics. The dimensions of the moving window are $12\leavevmode\nobreak\ $\mu\mathrm{m}$\times 12\leavevmode\nobreak\ $\mu\mathrm{m}$\times 20\leavevmode\nobreak\ $\mu\mathrm{m}$$, comprising $384\times 384\times 512$ cells in the $x$, $y$, and $z$ directions, respectively. Each cell contains 8 macro particles, which is sufficient to solve the electron dynamics effectively. For particle and field updates, the simulations utilise the Boris pusher [28] and the CKC Maxwell solver [29].
Figure 2: PIC results: SPs excitation on the flat CNT surface. (a) Density distribution of wall electrons (grey colourmap) and laser field (green-yellow colourmap). (b) Longitudinal acceleration field $E_{z}$. (c) Line plots of $E_{z}$ along $x=0\leavevmode\nobreak\ $\mu\mathrm{m}$$ (black) and $x=-2.5\leavevmode\nobreak\ $\mu\mathrm{m}$$ (green), indicated by the horizontal dashed lines in (b), respectively. (d) Line plots of $E_{z}$ along $z=46\leavevmode\nobreak\ $\mu\mathrm{m}$$ (orange), $z=48\leavevmode\nobreak\ $\mu\mathrm{m}$$ (green) and $z=49\leavevmode\nobreak\ $\mu\mathrm{m}$$ (black), indicated by the vertical dashed lines in (b) respectively. The vertical grey line indicates the position of the interface between the CNT wall and the vacuum. The corresponding theoretical estimations are shown by the dashed lines, respectively.
This physical phenomenon can be theoretically studied by solving Maxwell’s equations with appropriate boundary conditions at the surface [30]. The longitudinal electric field can be expressed as
$E_{z}=E_{z0}e^{-k_{x}x}e^{i(k_{z}z-\omega_{sp}t)}\mathrm{,}$ (3)
where $E_{z0}$ is the amplitude and $k_{x}=\omega_{sp}\sqrt{n^{2}-1}/c$. It is shown in Eq. (3) that the EM field decays exponentially away from the surface which can be observed from PIC simulations as shown in Fig. 2(c). Both longitudinal and transverse SP fields with an amplitude of $0.1\leavevmode\nobreak\ $\mathrm{T}\mathrm{V}\mathrm{/}\mathrm{m}$$ can leak into a small vacuum region near the surface and decay exponentially, as shown in Fig. 2(b) to (d). The resonance wavelength is theoretically calculated to be $\lambda_{sp,Theo}=3.24\leavevmode\nobreak\ $\mu\mathrm{m}$$, which agrees well with the simulation result $\lambda_{sp,PIC}\simeq 3.3\leavevmode\nobreak\ $\mu\mathrm{m}$$. SP decays slower in vacuum than in CNTs, since $k_{x,\text{vacuum}}=\omega_{sp}\sqrt{n^{2}-1}/c<k_{x,\text{CNTs}}=\omega_{sp}\sqrt{n^{2}-\varepsilon}/c$ [25]. The longitudinal component $E_{z}$ along the $x$-direction is plotted at three different positions by solid lines in Fig.2(d) and shows good agreement with the theoretical estimates indicated by the dashed lines. These agreements benchmark the validation of PIC simulations for SP dynamics. The observed deviations, e.g. at $z=50\leavevmode\nobreak\ $\mu\mathrm{m}$$ as shown by the solid and dashed black lines, are due to the influence of the laser field. With the evanescent nature of the leaky EM field from the flat surface, the accelerating volume is limited to a small region, e.g. $|\Delta x|<1/k_{sp}\leavevmode\nobreak\ \sim 0.2\leavevmode\nobreak\ $\mu\mathrm{m}$$ for our case, making the acceleration highly sensitive to the angular distribution of both the driving and witness beams. Consequently, the total charge and energy gain of the witness beam are limited. Therefore, it is not suitable for high-energy particle acceleration.
II.2 On cylindrical surface